OpenStudy (anonymous):

4 square root x -1=3

OpenStudy (popitree):

x=82

OpenStudy (popitree):

\[\sqrt[4]{x-1} = 3\]

jhonyy9 (jhonyy9):

4sqrt x - 1 = 3 4sqrt x = 4 sqrt x = 1 x =1